# On the convergence of adaptive approximations for stochastic differential equations

### v4

@@ 2.1 Rough path convergence of piecewise polynomial approximations
If $q=0$, then $\mathcal{W}_{n}=\{W_{t_{n+1}}-W_{t_{n}}\}$ and each $\widetilde{W}^{n}$ is the piecewise linear path that agrees with the Brownian motion $W$ at the discretization points $\{t_{n}\}$ given by the partition $\mathcal{D}_{n}$. More generally, $\widetilde{W}^{n}$ is a piecewise degree $q+1$ polynomial that matches “moment” information of $W$ on $\mathcal{D}_{n}$ (see [21, Theorem 2.4] for details).
@@
    For a fixed $q\geq 0$, the process $\big{\{}W_{t}-\mathbb{E}\big{[}W_{t}\hskip 0.7113pt|\hskip 0.7113pt\big{\{}\int_{0}^{1}s^{\hskip 0.7113ptp}\,dW_{s}\big{\}}_{0\hskip 0.7113pt\leq\hskip 0.7113ptp\hskip 0.7113pt\leq\hskip 0.7113ptq}\hskip 0.7113pt\big{]}\big{\}}_{t\in[0,1]}$ is a centered Gaussian process on $[0,1]$ and is independent of $\big{\{}\int_{0}^{1}s^{\hskip 0.7113ptp}\,dW_{s}\big{\}}_{0\hskip 0.7113pt\leq\hskip 0.7113ptp\hskip 0.7113pt\leq\hskip 0.7113ptq}$. This is a direct consequence of a polynomial expansion of Brownian motion (see, for example, [21, Theorem 2.2] or [26, Proposition 1.3] for more details).
@@
We shall now introduce the concepts of rough paths and $\alpha\hskip 0.7113pt$-Hölder regularity. These are the two key ingredients needed to establish SDE (or CDE) solutions as Lipschitz continuous functions of the Brownian motion (or driving path), see [22, 50].
@@
Now we have introduced some key definitions from rough path theory, we can strengthen the pointwise and mean squared convergence established in Theorem 2.11. To do so, we will largely follow the rough path analysis detailed in [23, Section 13.3.2].
@@
Let $\beta\in(\alpha,\frac{1}{2})$ denote a fixed constant. Then by [23, Corollary 13.14 (i)], there exists $c>0$ (depending on $\beta$ and $T$) such that
@@
Note that it immediately follows from the uniform Hölder bounds (43) and (44) that, almost surely, any subsequence of $\{(W-\widetilde{W}^{n},\mathbb{W}-\mathbb{\widetilde{W}}^{n})\}_{n\hskip 0.7113pt\geq\hskip 0.7113pt0}$ is uniformly bounded and uniformly equicontinuous in $(s,t)$. So by the Arzelà-Ascoli theorem, there will exist a further subsequence converging uniformly to some random continuous function $E:\{(s,t):0\leq s\leq t\leq T\}\rightarrow\mathbb{R}^{d}\oplus\mathbb{R}^{d\times d}$. Therefore, almost surely, for a fixed $(s,t)$, $E_{s,t}$ is the limit of a subsequence of $\{(W-\widetilde{W}^{n},\mathbb{W}-\mathbb{\widetilde{W}}^{n})\}$. However, by Theorems 2.11 and 2.14, the only possible limit is zero as $\boldsymbol{\widetilde{W}}{}_{\hskip-2.13394pts,t}^{n}$ converges to $\boldsymbol{W}_{\hskip-2.13394pts,t}$ almost surely. Since $E$ is uniformly continuous almost surely, it directly follows that it must be zero. Hence, almost surely, $\{(W-\widetilde{W}^{n},\mathbb{W}-\mathbb{\widetilde{W}}^{n})\}$ will converge uniformly to zero as $n\rightarrow\infty$.
@@ 2.2 Pathwise convergence of “polynomial-driven” CDEs with small local errors
Since we have shown $\big{\{}\boldsymbol{\widetilde{W}}^{n}\big{\}}$ converges to the Stratonovich Brownian motion $\boldsymbol{W}$ in the $\alpha\hskip 0.7113pt$-Hölder metric, we can apply an extension of the Universal Limit Theorem from rough path theory [22, Theorem 4.9] to show that the sequence of CDEs (where the $n$-th CDE is driven by $\widetilde{W}^{n}$) converges in the $\alpha\hskip 0.7113pt$-Hölder metric to the rough differential equation driven by $\boldsymbol{W}$ – which is just the Stratonovich SDE (1). To this end, we use the following definition to impose smoothness on the vector fields.
@@
The result follows from Theorem 2.18 along with the Lipschitz continuity of “Rough SDEs” with respect to the driving rough path, see [22, Theorem 4.9].
@@
Technically speaking, we define the rough path $\boldsymbol{y}=(1,y,\mathbbm{y})$ as the limit of $\{\boldsymbol{\widetilde{y}}^{\hskip 0.7113ptn}\}_{n\hskip 0.7113pt\geq\hskip 0.7113pt0}$ and, using [22, Theorem 4.9], $\boldsymbol{y}$ can be taken to be the solution of the rough differential equation driven by Stratonovich enhanced Brownian motion $\boldsymbol{W}$. In particular, from $\boldsymbol{y}$, we can define a process $y=\{y_{t}\}$ as $y_{t}=y_{0}+y_{0,t}$ for $t\in[0,T]$, which coincides almost surely with the usual solution of the Stratonovich SDE (1).
@@
To make this precise, we follow the error analysis presented in [24, Section 4.2]. However, as rough path theory was formulated after [24], we can use an extension of Davie’s estimate [23, Theorem 10.29] to establish a Lipschitz property of CDE flows.
@@
holds for some almost surely finite random variable $M$. The result directly follows by applying [22, Theorem 4.9] to the CDE solutions $\widetilde{x}^{\hskip 0.7113ptn}$ and $\widetilde{y}^{\hskip 0.7113ptn}$, with $\|\cdot\|_{m}=\|\cdot\|_{2}$.
Using the Lipschitz estimate (49), we now adapt the proof of [24, Theorem 4.3] to give our main result on pathwise convergence for adaptive approximations close to polynomial driven CDEs – such as Heun’s method (13) and the SRK method (14).
@@ 2.3 Numerical methods for SDEs viewed as approximations of CDEs
In [21], it was shown that the coefficients in the polynomial approximation $\widetilde{W}^{\hskip 0.7113ptn}$ can be expressed as linear combinations of rescaled iterated integrals with the form (57). In particular, this means that they can be uniformly bounded as $O(h_{k}^{\alpha})$ almost surely. For example, in the simplest cases where the polynomial’s degree is either $1$ or $2$, we have $\widetilde{W}_{t_{k}\hskip 0.7113pt,t}^{\hskip 0.7113ptn}=\frac{t-t_{k}}{h_{k}}\hskip 0.7113ptW_{k}$ or $\widetilde{W}_{t_{k}\hskip 0.7113pt,t}^{\hskip 0.7113ptn}=\frac{t-t_{k}}{h_{k}}\hskip 0.7113ptW_{k}+\frac{6(t_{k+1}-t)(t-t_{k})}{h_{k}^{2}}\hskip 0.7113ptH_{k}$ with $W_{k}\hskip 0.7113pt,H_{k}\sim O(h_{k}^{\alpha})$.
@@
Whilst a general formula for polynomial approximations of second iterated integrals (or equivalent Lévy areas) is given in [18, Theorem 5.4], we only use
@@ 3 Counterexample involving non-nested step sizes
We refer to [24, Section 4.1] for a counterexample with no-skip dyadic step sizes and the Euler-Maruyama method – which does not satisfy condition (3) in general.
@@ 4 Numerical example
In this section, we will consider the well-known SABR model used in mathematical finance [7, 11, 27, 47]. The model describes the evolution of an interest (or exchange) rate $S$ whose volatility $\sigma$ is also stochastic. It is given by
@@
For the non-previsible step size, we have modified the Proportional-Integral (PI) controller for commutative SDEs detailed in [6, 34] and implemented it in the popular ODE/SDE simulation package “Diffrax” [40]. This proposes a new step size $h_{k+1}$ as
@@
where $\{\hskip 0.7113pt\mathrm{Fac}_{\max}\hskip 0.7113pt,\mathrm{Fac}_{\min}\}$ are the maximum and minimum factors $h_{k}$ can change by, $\mathrm{Fac}\in(0,1)$ is a safety factor, $\{K_{I},K_{P}\}$ are the integral and proportional coefficients, $C>0$ is the absolute tolerance and $e(Y_{k},h_{k},\mathcal{W}_{k})$ denotes the rescaled local error estimator $\frac{1}{C\sqrt{d}}\|Y_{k+1}-\widetilde{Y}_{k+1}\|_{2}$ obtained by computing $(Y_{k+1}\hskip 0.7113pt,\widetilde{Y}_{k+1})$ from $(Y_{k},h_{k},\mathcal{W}_{k})$ [^3] In our case, the pair $(Y_{k+1}\hskip 0.7113pt,\widetilde{Y}_{k+1})$ will be obtained from an embedded Runge-Kutta method. Whilst it performed worse in our experiment, one can also compare one step and two half-steps of the same method. Of course, the numerical solution should then be propagated using the two half-steps.. If $e(Y_{k+1}\hskip 0.7113pt,h_{k+1}\hskip 0.7113pt,\mathcal{W}_{k+1})>1$, then the step $h_{k+1}$ is rejected and $\hskip 0.7113pt\mathrm{Fac}_{\max}$ is temporarily set equal to $\mathrm{Fac}<1$ so that the next candidate step size is strictly less than $h_{k}$. Otherwise, if the estimator satisfies $e(Y_{k+1}\hskip 0.7113pt,h_{k+1}\hskip 0.7113pt,\mathcal{W}_{k+1})\leq 1$, then $h_{k+1}$ is accepted. For PI adaptive SDE solvers, Diffrax always uses an initial candidate step size of $0.01$. We also set the factor parameters to their default values in Diffrax, $\mathrm{Fac}_{\max}=10$, $\mathrm{Fac}_{\min}=0.2,\mathrm{Fac}=0.9$ and, as recommended by [34], we choose $K_{I}=0.3,K_{P}=0.1$.
@@
A crucial aspect of the error estimator (65) is that $Y^{C}$ and $Y^{\text{fine}}$ are computed with respect to the same Brownian motion. Fortunately, this can be achieved in a convenient manner by Diffrax’s (single-seed) Virtual Brownian Tree [36]. However, we do note that (65) is an estimator for mean squared error – which is very different from the pathwise convergence established by our main result, Theorem 2.26.
@@ 5 Conclusion
    Examples of such schemes include high order weak approximations [37, 56, 57, 71] and the below extension of Heun’s method for general Itô SDEs introduced in [62]:
@@
  • Does the Lévy construction detailed in Appendix B extend to an explicit algorithm for generating $N$ terms in the polynomial expansion of Brownian motion? [21, 26]
    This would allow adaptive solvers to use accurate Lévy area approximations [18].
  • Do solvers that use non-previsible adaptive step sizes converge for stiff SDEs? [61]
@@ Appendix A Improved local accuracy of the SPaRK method (14)
It is also known that Heun’s method achieves second order weak convergence for SDEs with additive noise [46] (that is, when $g(y)=\sigma$ for a constant matrix $\sigma\in\mathbb{R}^{e\times d}$),
@@
Thus, by the standard mean squared error analysis of Milstein and Tretyakov [52], we see that stochastic Runge-Kutta method will converge strongly with rate $O(h^{\frac{3}{2}})$. Similarly, it is also clear that the method will achieve second order weak convergence. We refer the reader to [17] for more details on the analysis of commutative noise SDEs. For non-additive noise SDEs, it is straightforward to see that the proposed stochastic Runge-Kutta method exhibits the same strong convergence rates as Heun’s method.

### v5

@@ 2.1 Rough path convergence of piecewise polynomial approximations
If $q=0$, then $\mathcal{W}_{n}=\{W_{t_{n+1}}-W_{t_{n}}\}$ and each $\widetilde{W}^{n}$ is the piecewise linear path that agrees with the Brownian motion $W$ at the discretization points $\{t_{n}\}$ given by the partition $\mathcal{D}_{n}$. More generally, $\widetilde{W}^{n}$ is a piecewise degree $q+1$ polynomial that matches “moment” information of $W$ on $\mathcal{D}_{n}$ (see [foster2020poly, Theorem 2.4] for details).
@@
    For a fixed $q\geq 0$, the process $\big\{W_{t}-\mathbb{E}\big[W_{t}\hskip 0.7113pt|\hskip 0.7113pt\big\{\int_{0}^{1}s^{\hskip 0.7113ptp}\,dW_{s}\big\}_{0\hskip 0.7113pt\leq\hskip 0.7113ptp\hskip 0.7113pt\leq\hskip 0.7113ptq}\hskip 0.7113pt\big]\big\}_{t\in[0,1]}$ is a centered Gaussian process on $[0,1]$ and is independent of $\big\{\int_{0}^{1}s^{\hskip 0.7113ptp}\,dW_{s}\big\}_{0\hskip 0.7113pt\leq\hskip 0.7113ptp\hskip 0.7113pt\leq\hskip 0.7113ptq}$. This is a direct consequence of a polynomial expansion of Brownian motion (see, for example, [foster2020poly, Theorem 2.2] or [habermann2021poly, Proposition 1.3] for more details).
@@
We shall now introduce the concepts of rough paths and $\alpha\hskip 0.7113pt$-Hölder regularity. These are the two key ingredients needed to establish SDE (or CDE) solutions as Lipschitz continuous functions of the Brownian motion (or driving path), see [friz2024rsdes, lyons2007notes].
@@
Now we have introduced some key definitions from rough path theory, we can strengthen the pointwise and mean squared convergence established in Theorem 2.11. To do so, we will largely follow the rough path analysis detailed in [frizvictoir2010roughpaths, Section 13.3.2].
@@
Let $\beta\in(\alpha,\frac{1}{2})$ denote a fixed constant. Then by [frizvictoir2010roughpaths, Corollary 13.14 (i)], there exists $c>0$ (depending on $\beta$ and $T$) such that
@@
Note that it immediately follows from the uniform Hölder bounds (43) and (44) that, almost surely, any subsequence of $\{(W-\widetilde{W}^{n},\mathbb{W}-\mathbb{\widetilde{W}}^{n})\}_{n\hskip 0.7113pt\geq\hskip 0.7113pt0}$ is uniformly bounded and uniformly equicontinuous in $(s,t)$. So by the Arzelà-Ascoli theorem, there will exist a further subsequence converging uniformly to some random continuous function $E:\{(s,t):0\leq s\leq t\leq T\}\rightarrow\mathbb{R}^{d}\oplus\mathbb{R}^{d\times d}$. Therefore, almost surely, for a fixed $(s,t)$, $E_{s,t}$ is the limit of a subsequence of $\{(W-\widetilde{W}^{n},\mathbb{W}-\mathbb{\widetilde{W}}^{n})\}$. However, by Theorems 2.11 and 2.14, the only possible limit is zero as $\boldsymbol{\widetilde{W}}{}_{\hskip-2.13394pts,t}^{n}$ converges to $\boldsymbol{W}_{\hskip-2.13394pts,t}$ almost surely.
Since $E$ is uniformly continuous almost surely, it directly follows that it must be zero. Hence, almost surely, $\{(W-\widetilde{W}^{n},\mathbb{W}-\mathbb{\widetilde{W}}^{n})\}$ will converge uniformly to zero as $n\rightarrow\infty$.
@@ 2.2 Pathwise convergence of “polynomial-driven” CDEs with small local errors
Since we have shown $\big\{\boldsymbol{\widetilde{W}}^{n}\big\}$ converges to the Stratonovich Brownian motion $\boldsymbol{W}$ in the $\alpha\hskip 0.7113pt$-Hölder metric, we can apply an extension of the Universal Limit Theorem from rough path theory [friz2024rsdes, Theorem 4.9] to show that the sequence of CDEs (where the $n$-th CDE is driven by $\widetilde{W}^{n}$) converges in the $\alpha\hskip 0.7113pt$-Hölder metric to the rough differential equation driven by $\boldsymbol{W}$ – which is just the Stratonovich SDE (1). To this end, we use the following definition to impose smoothness on the vector fields.
@@
The result follows from Theorem 2.18 along with the Lipschitz continuity of “Rough SDEs” with respect to the driving rough path, see [friz2024rsdes, Theorem 4.9].
@@
Technically speaking, we define the rough path $\boldsymbol{y}=(1,y,\mathbbm{y})$ as the limit of $\{\boldsymbol{\widetilde{y}}^{\hskip 0.7113ptn}\}_{n\hskip 0.7113pt\geq\hskip 0.7113pt0}$ and, using [friz2024rsdes, Theorem 4.9], $\boldsymbol{y}$ can be taken to be the solution of the rough differential equation driven by Stratonovich enhanced Brownian motion $\boldsymbol{W}$. In particular, from $\boldsymbol{y}$, we can define a process $y=\{y_{t}\}$ as $y_{t}=y_{0}+y_{0,t}$ for $t\in[0,T]$, which coincides almost surely with the usual solution of the Stratonovich SDE (1).
@@
To make this precise, we follow the error analysis presented in [gaines1997variable, Section 4.2]. However, as rough path theory was formulated after [gaines1997variable], we can use an extension of Davie’s estimate [frizvictoir2010roughpaths, Theorem 10.29] to establish a Lipschitz property of CDE flows.
@@
holds for some almost surely finite random variable $M$. The result directly follows by applying [friz2024rsdes, Theorem 4.9] to the CDE solutions $\widetilde{x}^{\hskip 0.7113ptn}$ and $\widetilde{y}^{\hskip 0.7113ptn}$, with $\|\cdot\|_{m}=\|\cdot\|_{2}$.
Using the Lipschitz estimate (49), we now adapt the proof of [gaines1997variable, Theorem 4.3] to give our main result on pathwise convergence for adaptive approximations close to polynomial driven CDEs – such as Heun’s method (13) and the SRK method (14).
@@ 2.3 Numerical methods for SDEs viewed as approximations of CDEs
In [foster2020poly], it was shown that the coefficients in the polynomial approximation $\widetilde{W}^{\hskip 0.7113ptn}$ can be expressed as linear combinations of rescaled iterated integrals with the form (57). In particular, this means that they can be uniformly bounded as $O(h_{k}^{\alpha})$ almost surely. For example, in the simplest cases where the polynomial’s degree is either $1$ or $2$, we have $\widetilde{W}_{t_{k}\hskip 0.7113pt,t}^{\hskip 0.7113ptn}=\frac{t-t_{k}}{h_{k}}\hskip 0.7113ptW_{k}$ or $\widetilde{W}_{t_{k}\hskip 0.7113pt,t}^{\hskip 0.7113ptn}=\frac{t-t_{k}}{h_{k}}\hskip 0.7113ptW_{k}+\frac{6(t_{k+1}-t)(t-t_{k})}{h_{k}^{2}}\hskip 0.7113ptH_{k}$ with $W_{k}\hskip 0.7113pt,H_{k}\sim O(h_{k}^{\alpha})$.
@@
Whilst a general formula for polynomial approximations of second iterated integrals (or equivalent Lévy areas) is given in [foster2023levyarea, Theorem 5.4], we only use
@@ 3 Counterexample involving non-nested step sizes
We refer to [gaines1997variable, Section 4.1] for a counterexample with no-skip dyadic step sizes and the Euler-Maruyama method – which does not satisfy condition (3) in general.
@@ 4 Numerical example
In this section, we will consider the well-known SABR model used in mathematical finance [SABR2017Cai, SABR2018Cui, SABR2002Hagan, SABR2017Leitao]. The model describes the evolution of an interest (or exchange) rate $S$ whose volatility $\sigma$ is also stochastic. It is given by
@@
For the non-previsible step size, we have modified the Proportional-Integral (PI) controller for commutative SDEs detailed in [burrage2004variable, ilie2015adaptive] and implemented it in the popular ODE/SDE simulation package “Diffrax” [kidger2022nde]. This proposes a new step size $h_{k+1}$ as
@@
where $\{\hskip 0.7113pt\mathrm{Fac}_{\max}\hskip 0.7113pt,\mathrm{Fac}_{\min}\}$ are the maximum and minimum factors $h_{k}$ can change by, $\mathrm{Fac}\in(0,1)$ is a safety factor, $\{K_{I},K_{P}\}$ are the integral and proportional coefficients, $C>0$ is the absolute tolerance and $e(Y_{k},h_{k},\mathcal{W}_{k})$ denotes the rescaled local error estimator $\frac{1}{C\sqrt{d}}\|Y_{k+1}-\widetilde{Y}_{k+1}\|_{2}$ obtained by computing $(Y_{k+1}\hskip 0.7113pt,\widetilde{Y}_{k+1})$ from $(Y_{k},h_{k},\mathcal{W}_{k})$ [^3] In our case, the pair $(Y_{k+1}\hskip 0.7113pt,\widetilde{Y}_{k+1})$ will be obtained from an embedded Runge-Kutta method. Whilst it performed worse in our experiment, one can also compare one step and two half-steps of the same method. Of course, the numerical solution should then be propagated using the two half-steps.. If $e(Y_{k+1}\hskip 0.7113pt,h_{k+1}\hskip 0.7113pt,\mathcal{W}_{k+1})>1$, then the step $h_{k+1}$ is rejected and $\hskip 0.7113pt\mathrm{Fac}_{\max}$ is temporarily set equal to $\mathrm{Fac}<1$ so that the next candidate step size is strictly less than $h_{k}$. Otherwise, if the estimator satisfies $e(Y_{k+1}\hskip 0.7113pt,h_{k+1}\hskip 0.7113pt,\mathcal{W}_{k+1})\leq 1$, then $h_{k+1}$ is accepted. For PI adaptive SDE solvers, Diffrax always uses an initial candidate step size of $0.01$. We also set the factor parameters to their default values in Diffrax, $\mathrm{Fac}_{\max}=10$, $\mathrm{Fac}_{\min}=0.2,\mathrm{Fac}=0.9$ and, as recommended by [ilie2015adaptive], we choose $K_{I}=0.3,K_{P}=0.1$.
@@
A crucial aspect of the error estimator (65) is that $Y^{C}$ and $Y^{\text{fine}}$ are computed with respect to the same Brownian motion. Fortunately, this can be achieved in a convenient manner by Diffrax’s (single-seed) Virtual Brownian Tree [jelincic2024VBT]. However, we do note that (65) is an estimator for mean squared error – which is very different from the pathwise convergence established by our main result, Theorem 2.26.
@@ 5 Conclusion
    Examples of such schemes include high order weak approximations [jelincic2025levygan, ninomiya2009ninomiya, ninomiya2008victoir, talay1990HighOrder] and the below extension of Heun’s method for general Itô SDEs introduced in [roberts2012euler]:
@@
  • Does the Lévy construction detailed in Appendix B extend to an explicit algorithm for generating $N$ terms in the polynomial expansion of Brownian motion? [foster2020poly, habermann2021poly]
    This would allow adaptive solvers to use accurate Lévy area approximations [foster2023levyarea].
  • Do solvers that use non-previsible adaptive step sizes converge for stiff SDEs? [rackauckas2020stiff]
@@ Appendix A Improved local accuracy of the SPaRK method (14)
It is also known that Heun’s method achieves second order weak convergence for SDEs with additive noise [leimkuhler2014OLD] (that is, when $g(y)=\sigma$ for a constant matrix $\sigma\in\mathbb{R}^{e\times d}$),
@@
Thus, by the standard mean squared error analysis of Milstein and Tretyakov [milstein2004physics], we see that stochastic Runge-Kutta method will converge strongly with rate $O(h^{\frac{3}{2}})$. Similarly, it is also clear that the method will achieve second order weak convergence. We refer the reader to [foster2024splitting] for more details on the analysis of commutative noise SDEs. For non-additive noise SDEs, it is straightforward to see that the proposed stochastic Runge-Kutta method exhibits the same strong convergence rates as Heun’s method.